OpenStudy (pinknabastak):

PLEASE HELP ASAP!!! medals will be given! Solve: (if using calculator, make sure to write down what you are plugging in your calculator) write answers as fraction:

OpenStudy (pinknabastak):

\[-x+2y-4z-5u=27\] \[-2x-3y+3z-3u=9\] \[5x-4y+2z-3u=5\] \[-2x+4z-5u=29\]

OpenStudy (agent0smith):

This is a system of equations with 4 variables; you are going to want to use a calculator.

OpenStudy (pinknabastak):

haha ok.. i dont exactly have one on me right now...but i hope you can help with that lol

OpenStudy (agent0smith):

Plug them into wolfram alpha. I'll try it.

OpenStudy (agent0smith):

BTW you are almost guaranteed to need a calculator for this class. Buy one.

OpenStudy (pinknabastak):

i have one and i already took the class im just taking the final exam a year later haha...my calc is just not with me and i completely forgot how to do anything haha

OpenStudy (agent0smith):

You shouldn't try to do anything with solving 4 equations/4 variables by hand. Too much work. Three equations is bad enough.

OpenStudy (anonymous):

but its possible though, right?

OpenStudy (agent0smith):

Of course. Solving a million equations with a million variables by hand is possible, too. Probably not fun though.

OpenStudy (pinknabastak):

my paper says my answer has to be in fractions....and i kinda hav to show work or atleast what i plugged into the calc..

OpenStudy (agent0smith):

The answers aren't in fractions, so your paper is stupid. You'd write that you plugged the four equations into your equation solver, and pressed solve.

OpenStudy (pinknabastak):

LMFAO.. haha ok:)

OpenStudy (agent0smith):

I mean there's no way you "show work" from a calculator. You write in the equations, press solve. That's it.

OpenStudy (pinknabastak):

theres another one thats.... two equations its.....

OpenStudy (agent0smith):

Two is doable by hand.

OpenStudy (pinknabastak):

\[x-10y=24\] \[12y=x ^{2}+2x-48\]

OpenStudy (agent0smith):

Use x−10y=24 to get either x = or y = and then substitute that into the second equation

OpenStudy (agent0smith):

eg. x−10y=24 gives x = 24+10y replace every x in the second equation with (24+10y)

OpenStudy (anonymous):

u could use matrices or cramer's rule... @pinknabastak

OpenStudy (pinknabastak):

@genius12 which is.....?

OpenStudy (pinknabastak):

@agent0smith aso i put an equation in an equation...makes no sense

OpenStudy (anonymous):

do you what determinants are? @pinknabastak

OpenStudy (agent0smith):

lol it makes sense. It's called substitution. Since we have x = 24+10y replace every x in this \[\Large 12y=x ^{2}+2x-48\]

OpenStudy (pinknabastak):

haha nope:) @genius12

OpenStudy (pinknabastak):

so then what..??

OpenStudy (agent0smith):

Plug it in. Show me what you get.

OpenStudy (pinknabastak):

12y=(24+10y)^2+2(24+10y)-48

OpenStudy (agent0smith):

Good, now simplify/expand that all out, and solve for y.

OpenStudy (anonymous):

nothing @pinknabastak its unrelated to this...

OpenStudy (pinknabastak):

12y=48+40y

OpenStudy (pinknabastak):

-48=28y

OpenStudy (pinknabastak):

or -28y=48

OpenStudy (agent0smith):

\[\Large 12y=(24+10y)^2+2(24+10y)-48\]I'm not sure you expanded this correctly. You don't even have a y^2 term...

OpenStudy (pinknabastak):

haha yea did i mention im bad at this?

OpenStudy (agent0smith):

Expand it out slowly, bit by bit. Remember that (24+10y)^2 is the same as (24+10y)(24+10y)

OpenStudy (pinknabastak):

yea i did that

OpenStudy (agent0smith):

Then show your work.

OpenStudy (pinknabastak):

12y=24+10y+24+10y+48+20y-48 48's cross out combine y's and 24's 12y=48+40y

OpenStudy (agent0smith):

(24+10y)(24+10y) do this part. Remember to use FOIL or whatever you're used to.

OpenStudy (pinknabastak):

576 + 240y +240y + 100y

OpenStudy (pinknabastak):

576 + 580y

OpenStudy (agent0smith):

10y*10y is 100y^2. But how come you did it completely different to what you did above?

OpenStudy (agent0smith):

576 + 240y +240y + 100y^2

OpenStudy (pinknabastak):

ohh i forgot the y^2 part

OpenStudy (pinknabastak):

so its 576 + 480y + 100y^2

OpenStudy (agent0smith):

12y=24+10y+24+10y+48+20y-48 see the mistakes you made here? This is missing all that stuff from above.

OpenStudy (agent0smith):

Yes, now do this again\[\Large 12y=(24+10y)^2+2(24+10y)-48 \]

OpenStudy (pinknabastak):

12y= 576+480y+100y^2 +20y 12y= 100y^2 + 500y + 576

OpenStudy (agent0smith):

Yes, now solve for y using the quadratic formula .

OpenStudy (pinknabastak):

uh.... no calc here... and i forgot that long go haha

OpenStudy (agent0smith):

You're on a computer, so you have a calculator. Use google if necessary. Google the quadratic formula.

OpenStudy (pinknabastak):

\[\frac{ -500 \pm \sqrt{19,600} }{ 1,152 }\]

OpenStudy (agent0smith):

12y = 100y^2 + 500y + 576 Before using the quadratic formula, you have to get it equal to zero. You haven't used the quadratic formula correctly. It also shouldn't be something you forgot... memorize it. http://www.purplemath.com/modules/quadform.htm it's right near the top.

OpenStudy (pinknabastak):

its not working...

OpenStudy (agent0smith):

For ax2 + bx + c = 0, the value of x is given by: http://www.purplemath.com/modules/quads/qform01.gif

OpenStudy (pinknabastak):

i know thats the formula

OpenStudy (agent0smith):

Then... why'd you say you forgot it?

OpenStudy (nincompoop):

can you use matrices and other rules?

OpenStudy (nincompoop):

if not , you can ask for miracles and best hope to an answer

OpenStudy (agent0smith):

@nincompoop The current question is not the one at the top.

OpenStudy (nincompoop):

ermmmm laughing out loud nvm then

OpenStudy (pinknabastak):

i did forget it but i just looked it up and i already know..as in you didn tneed to put that last link.

OpenStudy (pinknabastak):

i got the vid to work!!!

OpenStudy (pinknabastak):

is that you in the vid?

OpenStudy (agent0smith):

Er what vid?

OpenStudy (anonymous):

SQUIRREL. ♥

OpenStudy (pinknabastak):

ok so now i got \[\frac{ 488 \pm \sqrt{7,744} }{ 200 }\]

OpenStudy (pinknabastak):

@Sally♡ huh?

OpenStudy (anonymous):

squiirrrreeeeeellllll ♥

OpenStudy (agent0smith):

Finish simplifying it.

OpenStudy (anonymous):

Whispers "squirrels are furry."

OpenStudy (anonymous):

Sally what are you talking about

OpenStudy (anonymous):

Sorry lol. I like squirrels (;

OpenStudy (agent0smith):

Also it's NEGATIVE b in the formula, you just have b.

OpenStudy (pinknabastak):

oh right..haha

OpenStudy (pinknabastak):

soo -400/200 or -576/200

OpenStudy (pinknabastak):

-2 or -2.88

OpenStudy (agent0smith):

Yes.

OpenStudy (pinknabastak):

and thats y

OpenStudy (pinknabastak):

wait but which one is it?

OpenStudy (agent0smith):

\[\Large x−10y=24\]\[\Large 12y=x^2+2x−48\] Now find x using the first equation. Use BOTH y's to get 2 diff x's

OpenStudy (pinknabastak):

i got 4 and 18.24

OpenStudy (agent0smith):

4 is right, 18.24 is not

OpenStudy (pinknabastak):

oops - 4.8

OpenStudy (agent0smith):

Yep.

OpenStudy (usukidoll):

O_O why not use matrices?

OpenStudy (pinknabastak):

ok so those two y's and x's are my answer?

OpenStudy (usukidoll):

I don't think you're done yet... you need to find x y z and u

OpenStudy (usukidoll):

this is an augmented matrix of 4 x 4

OpenStudy (pinknabastak):

its a different question half way through this chat i asked another question :)

OpenStudy (usukidoll):

oh *facepalm

OpenStudy (pinknabastak):

we just wolfram alpha for the top question...if you have another way oh doing the first one thoo id love to know it

OpenStudy (pinknabastak):

hello?? @UsukiDoll

OpenStudy (usukidoll):

4 x 4 matrix style

OpenStudy (usukidoll):